OpenStudy (vanessaperezxo):

MEDAL!!!! solve for x: 5|2x-2|+8=18 A- x= 2, x= -2 B- x= -2, x= 0 C- x= 2, x=0 D- x= 0, x= 1/2

OpenStudy (vanessaperezxo):

@horsegirl27

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

You cant have a negative as an answer because it will add to the |2x-2|. @vanessaperezxo

OpenStudy (anonymous):

C

OpenStudy (anonymous):

x=2 or x=0

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

Let's solve your equation step-by-step. 5(|2x−2|)+8=18 Step 1: Add -8 to both sides. 5(|2x−2|)+8+−8=18+−8 5(|2x−2|)=10 Step 2: Divide both sides by 5. 5(|2x−2|)5=105 |2x−2|=2 Step 3: Solve Absolute Value. |2x−2|=2 We know either 2x−2=2 or 2x−2=−2 2x−2=2(Possibility 1) 2x−2+2=2+2(Add 2 to both sides) 2x=4 2x2=42(Divide both sides by 2) x=2 2x−2=−2(Possibility 2) 2x−2+2=−2+2(Add 2 to both sides) 2x=0 2x2=02(Divide both sides by 2) x=0 Answer: x=2 or x=0

OpenStudy (anonymous):

Use this calculator. http://www.acalculator.com/math-calculators.html